# Uncovering Semantic Selectivity of Latent Groups in Higher Visual Cortex with Mutual Information-Guided Diffusion

### v1

@@ Abstract
Understanding how neural populations in higher visual areas encode object-centered visual information remains a central challenge in computational neuroscience. Prior works have investigated representational alignment between artificial neural networks and the visual cortex. Nevertheless, these findings are indirect and offer limited insights to the structure of neural populations themselves. Similarly, decoding-based methods have quantified semantic features from neural populations but have not uncovered their underlying organizations. This leaves open a scientific question: “how feature-specific visual information is distributed across neural populations in higher visual areas, and whether it is organized into structured, semantically meaningful subspaces.” To tackle this problem, we present MIG-Vis, a method that leverages the generative power of diffusion models to visualize and validate the visual-semantic attributes encoded in neural latent subspaces. Our method first uses a variational autoencoder to infer a group-wise disentangled neural latent subspace from neural populations. Subsequently, we propose a mutual information (MI)–guided diffusion synthesis procedure to visualize the specific visual-semantic features encoded by each latent group. We validate MIG-Vis on multi-session neural spiking datasets from the inferior temporal (IT) cortex of two macaques. The synthesized results demonstrate that our method identifies neural latent groups with clear semantic selectivity to diverse visual features, including object pose, inter-category transformations, and intra-class content. These findings provide direct, interpretable evidence of structured semantic representation in the higher visual cortex and advance our understanding of its encoding principles.
@@ 1 Introduction
Determining how populations of neurons in higher visual areas represent object‑centred visual information remains a central question in computational neuroscience (DiCarlo et al., 2012; Yamins & DiCarlo, 2016). A major line of research (Lindsey & Issa, 2024; Xie et al., 2024) has explored representational alignment between deep neural networks (DNNs) and primate visual cortex, showing that DNNs trained for object recognition—especially those with disentangled representation subspaces—closely resemble inferior temporal (IT) cortex activity. Nevertheless, these findings are indirect, relying on artificial model architecture and specially designed representational similarity metrics. Meanwhile, previous works have focused on single-unit selectivity (Rust & DiCarlo, 2010) or decoding-based methods (Freiwald & Tsao, 2010; Chang & Tsao, 2017) that quantify semantic features like object category or viewpoint in higher visual cortex. Earlier studies have also used pre-trained diffusion models with fMRI data (Luo et al., 2023a; b; Cerdas et al., 2024) to verify and classify that various brain regions specialize in processing certain category’s information.
However, existing approaches do not extract semantically interpretable neural representations from electrophysiological recordings in higher visual cortex. That said, no study has mapped the organization and structural patterns of higher visual area neural populations to distinct visual attributes. In practice, addressing this scientific problem is challenging due to neural single-unit activities in higher visual cortex exhibit mixed selectivity to multiple visual-semantic features (Chang & Tsao, 2017). We also verify this finding through an empirical study on the IT cortex of macaques in a passive object recognition task (Majaj et al., 2015) (the results are presented in Fig. 1).
To bridge this gap, we propose MIG‑Vis (short for Mutual Information-Guided Diffusion for Uncovering Semantic Selectivity of Neural Latent Groups in Higher Visual Cortex), a method identifying interpretable neural latent subspaces that exhibit visual-semantic selectivity. We propose to employ a disentangled latent variable model (LVM) to infer low-dimensional neural latents for semantic interpretation. We assume visual-semantic features span a multi-dimensional subspace encoding different aspects of object structure and variation. To capture this, we use a group-wise disentangled variational autoencoder (Li et al., 2025), which enforces each multi-dimensional latent group to represent distinct semantic features.
Given an identified latent group in the neural subspace, our subsequent goal is to characterize the specific visual-semantic features it encodes. Here we leverage the expressivity of generative diffusion models (Wang et al., 2023) to synthesize images guided by target neural signals (Luo et al., 2023a; Cerdas et al., 2024). We perform semantic image editing (Meng et al., 2021) with deterministic DDIM, which eliminates stochastic sampling noise to ensure that editing is completely guided by neural signals. This procedure enables us to visualize and interpret the specific visual-semantic features encoded by each neural latent group.
For diffusion guidance, prior works typically optimize simple statistical moments of target neural dimensions, such as absolute activation values (Luo et al., 2023a) or variance (Wang et al., 2024). Such approaches are especially common in the fMRI literature, where the neural space is directly manipulated—zero indicates no activation, and larger positive values correspond to stronger activation. However, such guidance strategies are ill-suited for uncovering semantic features of specific latent axes in our setting, since both positive and negative coordinates in the latent space can carry distinct meanings that first- or second-order moment objectives cannot fully capture.
To deal with this issue, we propose a novel guidance objective: maximizing mutual information (MI) between the synthesized image and the target latent group. MI provides an information-theoretic measure of inherent statistical dependence (Hjelm et al., 2018) between the synthesized image and the target neural latent group, capturing both linear and higher-order relationships. Therefore, maximizing MI guides the synthesized image to faithfully represent the full distribution of visual-semantic information in the target latent group, achieving alignment beyond simple first- or second-order moments.
@@
Figure 1: Single-neuron linear decoding results in IT cortex of two macaques during a passive object recognition task. (A) Decoding results on both two macaques reveal that IT neurons exhibit mixed selectivity. The connecting line thickness denotes the weight scale. These neurons contribute to both low-level pose attributes (e.g., rotation) and high-level semantic features (e.g., category id). (B) Images with variations along each visual-semantic feature.
@@
We experimentally evaluate the efficacy of MIG-Vis on multi-session neural spiking datasets (Majaj et al., 2015) from the IT cortex of two macaques during a passive object recognition task. Extensive diffusion-synthesized results demonstrate that MIG-Vis infers disentangled neural latent groups with visual-semantic selectivity. By synthesizing image with the guidance of maximizing the mutual information between the image output and each latent group, we characterize clear semantic specialization across the disentangled groups: inter-category variation, intra-category pose, and intra-category content details. Furthermore, we adopt this scientific interpretation method to within-group latent dimensions to uncover fine-grained semantic structure.
@@ 2 Preliminaries
Problem Formulation. For each single trial, data is composed of neural population and stimulus image pairs: $\left(\mathbf{x},\mathbf{y}\right)$. The neural population is denoted as $\mathbf{x}\in\mathbb{R}^{N}$, where $N$ is the number of recorded neural units. The corresponding image is represented as $\mathbf{y}\in\mathbb{R}^{1\times H\times W}$, where $H$ and $W$ denote the height and width of the gray-scale image. We develop a group-wise disentangled VAE based on Li et al. (2025) to infer the neural latent groups, denoted as $\mathbf{z}=\left[\mathbf{z}_{1},\ldots,\mathbf{z}_{G}\right]^{\top}\in\mathbb{R}^{D}$, in which $D$ is the latent dimension number and $G$ is the number of groups. Each latent group is set to have the same dimensionality $D_{g}$, thus $D=G\times D_{g}$. Our goal is to investigate the specific visual-semantic encoding of the $g$-th neural latent group $\mathbf{z}_{g}$.
Classifier-Guided Diffusion Models. Given a set of image data samples $\mathbf{y}$, a diffusion probabilistic model (Ho et al., 2020; Song et al., 2020b) estimates its density $p_{\text{data}}(\mathbf{y})$ by first perturbing the data points in a $T$-step forward process: $q\left(\mathbf{y}_{t}\mid\mathbf{y}_{0}\right):=\mathcal{N}\left(\mathbf{y}_{t};\sqrt{\alpha_{t}}\mathbf{y}_{0},\left(1-\alpha_{t}\right)\mathbf{I}\right)$, where $\{\alpha_{t}\}_{t=1}^{T}$ denotes a noise-schedule. In the reverse process, a neural network $\bm{\epsilon}_{\bm{\theta}}\left(\mathbf{y}_{t},t\right)$ parameterized by $\bm{\theta}$ is trained to estimate the introduced noise at each timestep $t$. On the other hand, classifier-guided diffusion models (Dhariwal & Nichol, 2021) enable conditional generation $p(\mathbf{y}\mid\mathbf{z})$ given a class label $\mathbf{z}$. Formally, given the class label $\mathbf{z}$ and a guidance scale $\gamma>0$, by Bayes’ rule, the conditional distribution of sampled images $\mathbf{y}$ would be: $p^{\gamma}(\mathbf{y}\mid\mathbf{z})\propto p(\mathbf{y})p(\mathbf{z}\mid\mathbf{y})^{\gamma}$. By taking the log-derivative on both sides, this gives:
$\nabla_{\mathbf{y}_{t}}\log p^{\gamma}\left(\mathbf{y}_{t}\mid\mathbf{z}\right)=\underbrace{\nabla_{\mathbf{y}_{t}}\log p_{\bm{\theta}}(\mathbf{y}_{t})}_{\text{ est. by }\bm{\epsilon}_{\theta}(\mathbf{y}_{t},t)}+\gamma\underbrace{\nabla_{\mathbf{y}_{t}}\log p_{\bm{\phi}}\left(\mathbf{z}\mid\mathbf{y}_{t}\right)}_{\text{est. by guidance}}.$ (1)
@@ 3 Methodology
Figure 2: (A) The semantic image editing procedure. This consists of a deterministic forward process from $t=0$ to an intermediate timestep $t=t^{\prime}$, followed by a neural-guided deterministic synthesis process back to $t=0$. (B) A schematic of the Mutual Information (MI) Landscape. This landscape is defined by our guidance objective, which maximizes the MI between a synthesized image and the neural latent group $\mathbf{z}_{1}$ of the target image (encoding object pose).
In the following, we first describe how MIG-Vis infers a neural latent subspace via a group-wise disentangled VAE. We then present our approach for visualizing and interpreting visual-semantic variations encoded by each latent group through mutual information-guided diffusion synthesis.
@@ 3.1 Inferring Group-wise Disentangled Neural Latent Subspace
To uncover interpretable visual-semantic structures in neural populations, we use a group-wise disentangled VAE that infers latent subspaces encoding distinct factors of variation. Standard disentangled VAEs (Higgins et al., 2017; Chen et al., 2018) assume each generative factor can be captured by a single independent latent dimension. However, this setting is limited for high-level visual attributes (e.g., category identity, 3D rotation), which require multi-dimensional latent codes.
Hence, we relax the single-dimension constraint by using a group‑wise disentangled VAE (Esmaeili et al., 2019; Li et al., 2025), which encourages statistical independency between multi-dimensional neural latent groups. As suggested by previous research works (Locatello et al., 2019), well-disentangled semantic latents seemingly cannot be inferred without (implicit) supervision. We incorporate certain image attributes (i.e., rotation angles and category identity) as priors to inform the latent subspace learning. The supervision labels are concatenated into a vector $\mathbf{u}\in\mathbb{R}^{M}$, where $M$ denotes its dimensionality. We decompose the neural latent vector $\mathbf{z}$ into supervised latent groups $\mathbf{z}^{(s)}$ and unsupervised latent groups $\mathbf{z}^{(u)}$, such that $\mathbf{z}=\left[\mathbf{z}^{(s)},\mathbf{z}^{(u)}\right]^{\top}$. The latent groups within $\mathbf{z}^{(s)}$ are informed by labels, while the groups in $\mathbf{z}^{(u)}$ are inferred without supervision. We propose to optimize the following lower bound of evidence $p_{\bm{\xi}}(\mathbf{x},\mathbf{u})$:
@@
where the probabilistic encoder $q_{\bm{\psi}}(\cdot)$ and decoder $p_{\bm{\xi}}(\cdot)$ are parameterized by $\bm{\psi}$ and $\bm{\xi}$, respectively. $q_{\bm{\psi}}(\mathbf{z})=\sum_{n=1}^{N}q_{\bm{\psi}}\left(\mathbf{z}\mid\mathbf{x}^{(n)}\right)q\left(\mathbf{x}^{(n)}\right)$ is the aggregated posterior over a sample set of size $N$, $g\in\{1,2,\ldots,G\}$ denotes the latent group index, and hyperparameter $\beta$ controls the penalty scale. We note that the group-wise factorized density in the partial correlation (PC) term is intractable, thus we use the importance sampling (IS) estimator (Li et al., 2025) to approximate the PC during training. Importantly, the use of weak-supervision labels and the PC penalty here does not degrade the neural reconstruction quality, as verified by the results in Section 4.
Given an inferred neural latent $\mathbf{z}$, the subsequent target is to characterize the specific visual-semantic factors encoded by a target latent group $\mathbf{z}_{g}$ within it. The efficacy of traditional latent traversal with neural decoders is limited in practice, as latent dimensions within a group may differ in scale, and simple perturbations to them often result in low-fidelity or uninterpretable images. We present empirical evaluations of these methods in Section 4.
@@ 3.2 Mutual Information Maximization Guidance
To tackle this challenge, we use a classifier-guided diffusion (Eq. 1). We adopt it due to its explicit conditional gradient term provides scientific interpretability. Our goal is to approximate the two terms on the RHS of Eq. 1. The unconditional score term is estimated by a denoiser $\bm{\epsilon}_{\theta}(\mathbf{y}_{t},t)$. To faithfully capture the semantic features encoded in latent group $\mathbf{z}_{g}$, we propose an MI-based guidance objective to construct the classifier. From it, we derive the conditional score $\nabla_{\mathbf{y}_{t}}\log p\left(\mathbf{z}\mid\mathbf{y}_{t}\right)$.
Mutual Information-Maximization Guidance Objective. Our goal is to characterize the visual-semantic features encoded in each neural latent group, capturing the entire statistical information rather than just first- or second-order statistics. To this end, we propose the following workflow: given an original image $\mathbf{y}$ and a target latent group $\mathbf{z}_{g}$, we first perturb the original image and then denoise it with the diffusion model by guiding the output to maximize its group-wise mutual information with $\mathbf{z}_{g}$. The MI between $\mathbf{z}_{g}$ and $\mathbf{y}$ is given by:
@@
Note that the MI captures the entire statistical dependence between $\mathbf{z}_{g}$ and the synthesized image $\mathbf{y}$. Given the fixed $\mathbf{z}_{g}$, the guidance objective of maximizing the MI enforces the synthesized $\mathbf{y}$ to faithfully contain the semantic information encoded in $\mathbf{z}_{g}$. Based on Eq. 1, our classifier-guided conditional score is:
$\nabla_{\mathbf{y}_{t}}\log p^{\gamma}\left(\mathbf{y}_{t}\mid\mathbf{z}\right)=\nabla_{\mathbf{y}_{t}}\log p_{\bm{\theta}}(\mathbf{y}_{t})+\gamma\nabla_{\mathbf{y}_{t}}\mathbf{MI}\left(\mathbf{z}_{g},\mathbf{y}_{t}\right).$ (4)
@@
During MI-guided image synthesis, at each step $t$, we first estimate its noise‑free approximation by $\hat{\mathbf{y}}_{0}=\left(\mathbf{y}_{t}-\sqrt{1-\alpha_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{y}_{t},t)\right)/\sqrt{\alpha_{t}}$, which is then used as input to the $s_{\bm{\phi}}\left(\cdot\right)$ network. Figure 2(B) illustrates our MI guidance objective, which guides image synthesis toward the first latent group $\mathbf{z}_{1}$ of the target image, with its strength controlled by the guidance scale $\gamma$.

### v2

@@ Abstract
Understanding how neural populations in higher visual areas encode object-centered visual information remains a key challenge in computational neuroscience. Prior work has examined representational alignment between artificial neural networks and the visual cortex, but such findings are indirect and provide limited insight into the structure of the neural population coding. Decoding-based methods can recover semantic features from neural activity, yet they do not reveal how these features are organized. This leaves an open question: how feature-specific visual information is distributed across neural populations, and whether it forms structured, semantically meaningful subspaces. To address this, we introduce MIG-Vis, a method that leverages diffusion models to visualize and validate visual-semantic attributes encoded in neural latent subspaces. MIG-Vis first learns a group-wise disentangled neural latent representation using a variational autoencoder. It then uses mutual information (MI)–guided diffusion synthesis to visualize the visual-semantic features encoded in each latent group. We validate MIG-Vis on multi-session neural spiking datasets from the inferior temporal (IT) cortex of two macaques. The synthesized results show that MIG-Vis identifies neural latent groups with clear semantic selectivity to diverse visual features, including object pose, inter-category variation, and intra-category content. These findings provide direct and interpretable evidence of structured semantic representation in higher visual cortex. The code repository of MIG-Vis is available at: https://github.com/BRAINML-GT/MIG-Vis.
@@ 1 Introduction
Determining how populations of neurons in higher visual areas represent object‑centred visual information remains a central question in computational neuroscience (DiCarlo et al., 2012; Yamins and DiCarlo, 2016). A major line of research (Lindsey and Issa, 2024; Xie et al., 2024) has explored representational alignment between deep neural networks (DNNs) and primate visual cortex, showing that DNNs trained for object recognition—especially those with disentangled representation subspaces—closely resemble inferior temporal (IT) cortex activity. Nevertheless, these findings are indirect, relying on artificial model architecture and specially designed representational similarity metrics. Meanwhile, previous works have focused on single-unit selectivity (Rust and DiCarlo, 2010) or decoding-based methods (Freiwald and Tsao, 2010; Chang and Tsao, 2017) that quantify semantic features like object category or viewpoint in higher visual cortex. Earlier studies have also used pre-trained diffusion models with fMRI data (Luo et al., 2023b; a; Cerdas et al., 2024) to verify and classify that various brain regions specialize in processing certain categories’ information.
@@
Figure 1: (A) Single-neuron decoding results show that IT neurons exhibit mixed selectivity, contributing to both low-level pose attributes (e.g., rotation) and high-level semantic features (e.g., category identity) to different extents (indicated by line thickness). (B) Example images varying along each visual-semantic feature.
However, existing approaches do not extract semantically interpretable neural representations from electrophysiological recordings in higher visual cortex. No study has mapped the organization and structural patterns of higher visual area neural populations to distinct visual attributes. In practice, addressing this scientific problem is challenging due to single-unit neural activities in the higher visual cortex exhibit mixed selectivity to multiple visual-semantic features (Chang and Tsao, 2017). We also verify this finding through an empirical study on the IT cortex of macaques in a passive object recognition task (Majaj et al., 2015) (the results are presented in Fig. 1).
To address this gap, we propose MIG-Vis (Mutual Information-Guided Diffusion for uncovering semantic selectivity of neural latent groups in higher Visual cortex), a method that identifies interpretable neural latent subspaces and visualizes the semantic meaning encoded in each subspace. We assume that multiple latent dimensions form a group that encodes a specific type of semantic feature. For example, object category may be represented by one latent group, while rotation is represented by another. To capture this structure, we use a group-wise disentangled variational autoencoder (Li et al., 2025), which learns latent groups composed of multiple dimensions, each corresponding to a distinct semantic feature type. Within each group, individual dimensions can still capture different aspects of that feature type, such as shape, texture, or lighting within a category-related group.
Given the learned neural latent space, our next goal is to understand the visual-semantic features that a certain latent group encodes. We achieve this by perturbing the latent and generating corresponding images, then comparing them to the original image to observe the semantic changes introduced by the perturbation. Traditionally, this can be done using a neural-to-image decoder to map a perturbed neural latent to an image (Whiteway et al., 2021). However, such a decoder tends to produce a single best reconstruction and may smooth out subtle variations in the latent space. As a result, small perturbations may not lead to clearly distinguishable semantic changes. An alternative is to use diffusion with guidance. Prior work typically guides diffusion by optimizing simple statistical moments of target neural dimensions, such as absolute activation values (Luo et al., 2023b) or variance (Wang et al., 2024). Such approaches are especially common in the fMRI literature, where neural activity is directly manipulated: zero indicates no activation, and larger positive values correspond to stronger responses. In our case, the neural representation lies in a learned latent space where both positive and negative values carry meaningful but distinct semantics. Thus, maximizing magnitude or variance does not result in meaningful semantic change encoded in neural latent space.
To address limitations in existing methods, we perturb the latent representation by adding or subtracting values along specific dimensions and seek to generate images that reflect these perturbed latents. Rather than relying on a decoder or diffusion guided by magnitude or variance maximization, we steer the diffusion process by maximizing the mutual information (MI) (Hjelm et al., 2018) between the synthesized image and the perturbed latent. MI captures the full statistical dependence between them, encouraging the generated image to retain the semantic changes introduced by the perturbation. Compared to direct neural-to-image decoding, MI-guided diffusion reduces the risk of collapsing semantic variations into an averaged reconstruction and is better suited to reveal how visual meaning changes along different latent directions.
@@
We evaluate MIG-Vis on multi-session neural spiking datasets (Majaj et al., 2015) from the IT cortex of two macaques performing a passive object recognition task. Diffusion-based synthesis shows that MIG-Vis learns disentangled neural latent groups with distinct visual-semantic selectivity. By generating images guided by maximizing mutual information between the output and each latent group, we identify clear semantic roles across groups, including pose, inter-category variation, and intra-category content details. We further apply this interpretation method within each group to reveal fine-grained semantic structure.
@@ 2 Preliminaries
Problem Formulation. For each single trial, data is composed of neural population and stimulus image pairs: $\left(\mathbf{x},\mathbf{y}\right)$. The neural population is denoted as $\mathbf{x}\in\mathbb{R}^{N}$, where $N$ is the number of recorded neural units. The corresponding image is represented as $\mathbf{y}\in\mathbb{R}^{1\times H\times W}$, where $H$ and $W$ denote the height and width of the gray-scale image. We develop a group-wise disentangled VAE based on Li et al. (2025) to infer the neural latent groups, denoted as $\mathbf{z}=\left[\mathbf{z}_{1},\ldots,\mathbf{z}_{G}\right]^{\top}\in\mathbb{R}^{D}$, in which $D$ is the latent dimension number and $G$ is the number of groups. Each latent group is set to have the same dimensionality $d_{g}$, thus $D=G\times d_{g}$. Our ultimate goal is to investigate the specific visual-semantic encoding of the $g$-th neural latent group.
Classifier-Guided Diffusion Models. Given a set of image data samples $\mathbf{y}$, a diffusion probabilistic model (Ho et al., 2020; Wang et al., 2023) estimates its density $p_{\text{data}}(\mathbf{y})$ by first perturbing the data points in a $T$-step forward process: $q\left(\mathbf{y}_{t}\mid\mathbf{y}_{0}\right):=\mathcal{N}\left(\mathbf{y}_{t};\sqrt{\alpha_{t}}\mathbf{y}_{0},\left(1-\alpha_{t}\right)\mathbf{I}\right)$, where $\{\alpha_{t}\}_{t=1}^{T}$ denotes a noise-schedule. In the reverse process, a neural network $\bm{\epsilon}_{\bm{\theta}}\left(\mathbf{y}_{t},t\right)$ parameterized by $\bm{\theta}$ is trained to estimate the introduced noise at each timestep $t$. On the other hand, classifier-guided diffusion models (Dhariwal and Nichol, 2021) enable conditional generation $p(\mathbf{y}\mid\mathbf{z})$ given a class label $\mathbf{z}$. Formally, given the class label $\mathbf{z}$ and a guidance scale $\eta>0$, by Bayes’ rule, the conditional distribution of sampled images $\mathbf{y}$ would be: $p^{\eta}(\mathbf{y}\mid\mathbf{z})\propto p(\mathbf{y})p(\mathbf{z}\mid\mathbf{y})^{\eta}$. By taking the log-derivative on both sides, this gives:
$\nabla_{\mathbf{y}_{t}}\log p^{\eta}\left(\mathbf{y}_{t}\mid\mathbf{z}\right)=\underbrace{\nabla_{\mathbf{y}_{t}}\log p_{\bm{\theta}}(\mathbf{y}_{t})}_{\text{ est. by }\bm{\epsilon}_{\theta}(\mathbf{y}_{t},t)}+\eta\underbrace{\nabla_{\mathbf{y}_{t}}\log p_{\bm{\phi}}\left(\mathbf{z}\mid\mathbf{y}_{t}\right)}_{\text{est. by guidance}}.$ (1)
@@ 3 Methodology
Figure 2: (A) The semantic image editing procedure. This consists of a deterministic forward process from $t=0$ to an intermediate timestep $t=t^{\prime}$, followed by a neural-guided deterministic synthesis process back to $t=0$. (B) Perturbation process. We show the neural manifold of the latent group $\mathbf{z}_{1}$, which encodes object pose. We traverse along its first group. Starting from the original image, we slightly adjust the latent variable to obtain perturbed latents and generate the corresponding images by maximizing mutual information with these perturbed representations. Since the perturbations are applied within the pose subspace, the resulting image changes consistently reflect rotation variations.
In the following, we first describe how MIG-Vis infers a neural latent space via a group-wise disentangled VAE. We then present our approach for visualizing and interpreting visual-semantic variations encoded by each latent group through mutual information-guided diffusion synthesis.
@@ 3.1 Inferring Group-wise Disentangled Neural Latent Subspace
To uncover interpretable visual–semantic structure in neural populations, we use a group-wise disentangled VAE that learns latent subspaces corresponding to different semantic factors. Traditional disentangled VAEs (Higgins et al., 2017; Chen et al., 2018) assume that each semantic factor is represented by a single independent latent dimension. However, this assumption is restrictive for high-level visual attributes, such as object category or 3D rotation, which typically require multiple latent dimensions to be represented adequately.
Hence, we relax the single-dimension assumption by using a group‑wise disentangled VAE (Esmaeili et al., 2019; Li et al., 2025), which encourages statistical independence between multi-dimensional neural latent groups. As suggested by previous works (Locatello et al., 2019), well-disentangled semantic latents seemingly cannot be inferred without (implicit) supervision. We incorporate certain image attributes (i.e., rotation angles and category identity) as supervision to inform the latent subspace learning. The supervision labels are concatenated into a vector $\mathbf{u}\in\mathbb{R}^{M}$, where $M$ denotes its dimensionality. We decompose the neural latent vector $\mathbf{z}$ into supervised latent groups $\mathbf{z}^{(s)}$ and unsupervised latent groups $\mathbf{z}^{(u)}$, such that $\mathbf{z}=\left[\mathbf{z}^{(s)},\mathbf{z}^{(u)}\right]^{\top}$. The latent groups within $\mathbf{z}^{(s)}$ are informed by labels, while the groups in $\mathbf{z}^{(u)}$ are inferred without supervision. We propose to optimize the following lower bound of evidence $p_{\bm{\xi}}(\mathbf{x},\mathbf{u})$:
@@
where the probabilistic encoder $q_{\bm{\psi}}(\cdot)$ and decoder $p_{\bm{\xi}}(\cdot)$ are parameterized by $\bm{\psi}$ and $\bm{\xi}$, respectively. $q_{\bm{\psi}}(\mathbf{z})=\sum_{n=1}^{N}q_{\bm{\psi}}\left(\mathbf{z}\mid\mathbf{x}^{(n)}\right)q\left(\mathbf{x}^{(n)}\right)$ is the aggregated posterior over a sample set of size $N$, $g\in\{1,2,\ldots,G\}$ denotes the latent group index, and hyperparameter $\beta$ controls the penalty scale. We note that the group-wise factorized density in the partial correlation (PC) term is intractable, thus we use the importance sampling (IS) estimator (Li et al., 2023; 2025) to approximate the PC during training. Importantly, the use of weak-supervision labels and the PC penalty here does not degrade the neural reconstruction quality, as verified by the results in Section 4.
@@ 3.2 Mutual Information Maximization Guidance
Given an inferred neural latent $\mathbf{z}$, the next goal is to characterize the visual–semantic factors encoded by a specific latent group $\mathbf{z}_{g}$. By perturbing the neural latent $\mathbf{z}_{g}$, we observe the resulting changes in the generated images, which helps reveal the underlying semantic variations. Traditionally, this is done by learning a neural-to-image decoder that directly maps the perturbed latent to an image. However, such a decoder often captures only the most dominant patterns and may miss subtle but meaningful variations in the latent. Instead, we use diffusion to generate an image that maximizes its mutual information with the perturbed latent. This encourages the synthesized image to preserve all relevant information encoded in the latent, rather than averaging it into a single reconstruction. As a result, the generated image better reflects the semantic changes introduced by the perturbation. In this sense, maximizing mutual information provides a more informative guidance signal than directly mapping from $\mathbf{z}$ to $\mathbf{y}$ using a standard decoder.
Mutual Information-Maximization Guidance. We start with an original image $\mathbf{y}$ and its corresponding neural latent group $\mathbf{z}_{g}$. To visualize the semantic features encoded in $\mathbf{z}_{g}$, we perturb it as $\tilde{\mathbf{z}}_{g}=\mathbf{z}_{g}+\gamma\mathbf{1};\gamma\sim\mathcal{U}(-a,a)$, where $a$ is a predefined half-range and $\mathbf{1}\in\mathbb{R}^{d_{g}}$ is an all-ones vector matching the dimension of the $g$-th latent group. This perturbation moves $\mathbf{z}_{g}$ in both positive and negative directions. We then add noise to the image $\mathbf{y}$ and use the diffusion model to gradually denoise it, while guiding the process so that the generated image stays closely related to $\tilde{\mathbf{z}}_{g}$. Because $\tilde{\mathbf{z}}_{g}$ is slightly perturbed from the original latent $\mathbf{z}_{g}$, the new image $\tilde{\mathbf{y}}$ is also a modified version of $\mathbf{y}$. The difference between the two images shows the semantic change caused by moving in the neural latent space. Concretely, we use classifier-guided diffusion (Eq. 1), since its explicit conditional gradient term makes scientific interpretation more accessible. We aim to approximate the two terms on the RHS of Eq. 1. The unconditional score term is estimated using a denoiser $\bm{\epsilon}_{\theta}(\mathbf{y}_{t},t)$. We then introduce an MI-based guidance to construct the classifier, from which we derive the conditional score $\nabla_{\mathbf{y}_{t}}\log p\left(\mathbf{z}\mid\mathbf{y}_{t}\right)$. Formally, the MI between $\mathbf{z}_{g}$ and $\mathbf{y}$ is given by:
@@
Note that mutual information captures the full statistical dependence between $\mathbf{z}_{g}$ and the image $\mathbf{y}$. During diffusion sampling, given the perturbed $\tilde{\mathbf{z}}_{g}$, maximizing mutual information guides the process so that the synthesized image $\tilde{\mathbf{y}}$ reflects the semantic information encoded in $\tilde{\mathbf{z}}_{g}$. Based on Eq. 1, the classifier-guided conditional score becomes:
$\nabla_{\mathbf{y}_{t}}\log p^{\eta}\left(\mathbf{y}_{t}\mid\mathbf{z}_{g}\right)=\nabla_{\mathbf{y}_{t}}\log p_{\bm{\theta}}(\mathbf{y}_{t})+\eta\nabla_{\mathbf{y}_{t}}\mathbf{MI}\left(\mathbf{z}_{g},\mathbf{y}_{t}\right).$ (4)
@@
During MI-guided image synthesis, at each step $t$, we first estimate its noise‑free approximation by $\hat{\mathbf{y}}_{0}=\left(\mathbf{y}_{t}-\sqrt{1-\alpha_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{y}_{t},t)\right)/\sqrt{\alpha_{t}}$, which is then used as input to the $s_{\bm{\phi}}\left(\cdot\right)$ network. Fig. 2(B) shows the perturbation of the neural latent and the corresponding images generated under MI guidance. MI steers image synthesis within the first latent group $\mathbf{z}_{1}$, with the guidance strength controlled by the scale parameter $\gamma$.